OpenStudy (anonymous):

Evaluate if x = -4 and y = 13. 3(2x -y) -63 -15 82 90

OpenStudy (anonymous):

plug in the numbers where u see x and distribute the 3

OpenStudy (anonymous):

I did and My answer wasnt a choice :(

OpenStudy (anonymous):

-24-26 = -50

OpenStudy (anonymous):

so if x = -4.. 3(2(-4) - y) and y is 13 so .. 3(2(-4) - 13) right?

OpenStudy (anonymous):

so 3(-8 - 13) and distribute = 3(-8) - 3(13) =

OpenStudy (anonymous):

-50

OpenStudy (anonymous):

13(3) is not 26

OpenStudy (anonymous):

Ok everyone just like is totally confusing me haha

OpenStudy (anonymous):

stick with me :) 3(-8) - 3(13) =

OpenStudy (anonymous):

oh yeah nameless your right i thought it was times 2

OpenStudy (anonymous):

-24 - 39 = -24 + -39 = -63

OpenStudy (anonymous):

-24-39 = -63

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

Thanks :D